# Decoupling Local Primordial non-Gaussianity from Relativistic Effects in the Galaxy Bispectrum

### v2

@@ Appendix D MegaMapper LBG biases
[…]
@@
Eq. D.3 is the K-correction which accounts for the redshifting effect on galaxy fluxes which are measured in fixed wavelength bands. $m_{\rm UV}-m$ is negligible for the MegaMapper LBG sample, so we choose to ignore it. $D_{\rm L}(z)$ is the luminosity distance and the best-fit parameters for equation Eq. D.1 are given by Table 3 of [83, reported also here in Table 3]. This paper gives a full description of the properties of LBGs and their selection process.
$z_{\text{eff }}$ $M_{\rm UV}^{\star}$ $\phi^{*}/\left(10\,h^{-1}\,\mathrm{Mpc}\right)^{-3}$ $\alpha$ $m_{\rm UV}^{\star}$
$2.0$ $-20.60$ $9.70$ $-1.60$ $24.2$
$3.0$ $-20.86$ $5.04$ $-1.78$ $24.7$
$3.8$ $-20.63$ $9.25$ $-1.57$ $25.4$
$4.9$ $-20.96$ $3.22$ $-1.60$ $25.5$
$5.9$ $-20.91$ $1.64$ $-1.87$ $25.8$
Table 3: Luminosity function best-fit parameters as given by [83], which gives references for deriving the best-fit values at each redshift.
@@
$n_{\mathrm{g}}\left(z,M_{\mathrm{c}}\right)=\int_{-\infty}^{M_{\mathrm{c}}(z)}\,\mathrm{d}M\,\varPhi(z,M),$ (D.4)
@@
$\mathcal{E}=-\frac{\mathrm{d}\ln n_{\mathrm{g}}}{\mathrm{d}\ln(1+z)}-2\,\left[1+\frac{1}{r\,{\cal H}}+\frac{2\,\ln 10}{5}\,\frac{\mathrm{d}K}{\mathrm{d}\ln(1+z)}\right]\mathcal{Q}\;.$ (D.6)
@@
$A(m)=-0.98(m-25)+0.11\;,\,\,\,\,\,\,\,\,B(m)=0.12(m-25)+0.17\;,$ (D.8)
which yields Eq. A.9 for a magnitude cut $m=24.5\,$.
@@
(b) Comoving number density $n_{g}$
@@
Figure 9: MegaMapper LBG survey specifications derived from [83, 80].
@@ D.1 Luminosity derivatives
Figure 10: Comparison of the shift on the value of $f_{\mathrm{NL}}$ induced by negelectig GR corrections for MegaMapper without luminosity derivatives (solid line, as shown in Fig. 7) and with luminosity derivatives (diamond).
@@ Appendix E Survey Parameters
Table 4: Survey parameters for Roman-like luminosity Model 1 ($F_{c}=1\times 10^{-16}\mathrm{erg}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$).
[…]
$0.5$ $0.25$ $9.33$ $-4.91$ $1.07$
$0.6$ $0.32$ $7.82$ $-4.71$ $1.14$
$0.7$ $0.39$ $6.65$ $-4.53$ $1.21$
$0.8$ $0.45$ $5.70$ $-4.38$ $1.29$
$0.9$ $0.51$ $4.90$ $-4.24$ $1.36$
$1.0$ $0.56$ $4.22$ $-4.12$ $1.44$
$1.1$ $0.61$ $3.64$ $-4.02$ $1.52$
$1.2$ $0.65$ $3.13$ $-3.92$ $1.60$
$1.3$ $0.68$ $2.69$ $-3.84$ $1.69$
$1.4$ $0.71$ $2.31$ $-3.76$ $1.77$
$1.5$ $0.74$ $1.97$ $-3.68$ $1.85$
$1.6$ $0.76$ $1.69$ $-3.60$ $1.93$
$1.7$ $0.78$ $1.44$ $-3.53$ $2.01$
$1.8$ $0.79$ $1.23$ $-3.45$ $2.09$
$1.9$ $0.80$ $1.05$ $-3.37$ $2.16$
@@
Table 5: Survey parameters for Roman-like luminosity Model 3 ($F_{c}=1\times 10^{-16}\mathrm{erg}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$).
[…]
$0.1$ $0.11$ $38.38$ $-2.66$ $0.55$
$0.2$ $0.38$ $15.88$ $-3.16$ $0.94$
$0.3$ $0.76$ $6.42$ $-3.83$ $1.50$
$0.4$ $1.21$ $2.17$ $-4.75$ $2.31$
$0.5$ $1.69$ $0.56$ $-6.03$ $3.44$
$0.6$ $2.16$ $0.10$ $-7.76$ $4.94$
Table 6: DESI BGS survey parameters ($m_{c}=20$).
[…]
$0.7$ $2.78$ $2.72$ $-4.33$ $1.66$
$0.8$ $3.23$ $1.98$ $-4.76$ $1.87$
$0.9$ $3.65$ $1.46$ $-5.19$ $2.09$
$1.0$ $4.03$ $1.09$ $-5.62$ $2.30$
$1.1$ $4.36$ $0.82$ $-6.05$ $2.52$
$1.2$ $4.65$ $0.62$ $-6.48$ $2.73$
$1.3$ $4.90$ $0.48$ $-4.91$ $2.94$
$1.4$ $5.12$ $0.34$ $-5.33$ $3.15$
$1.5$ $5.30$ $0.25$ $-5.74$ $3.35$
$1.6$ $5.45$ $0.18$ $-6.15$ $3.56$
$1.7$ $5.58$ $0.13$ $-6.55$ $3.76$
$1.8$ $5.68$ $0.10$ $-6.94$ $3.95$
$1.9$ $5.76$ $0.07$ $-7.32$ $4.14$
$2.0$ $5.83$ $0.06$ $-7.69$ $4.33$
@@
Table 7: Survey parameters for Euclid-like luminosity Model 1, benchmark flux cut ($F_{c}=3\times 10^{-16}\mathrm{erg}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$).
[…]
$0.9$ $3.65$ $1.56$ $-6.13$ $1.97$
$1.0$ $4.03$ $1.26$ $-5.93$ $2.07$
$1.1$ $4.36$ $1.02$ $-5.73$ $2.17$
$1.2$ $4.65$ $0.82$ $-5.53$ $2.26$
$1.3$ $4.90$ $0.67$ $-5.32$ $2.34$
$1.4$ $5.12$ $0.54$ $-5.11$ $2.41$
$1.5$ $5.30$ $0.44$ $-4.91$ $2.47$
$1.6$ $5.45$ $0.36$ $-4.71$ $2.52$
$1.7$ $5.58$ $0.29$ $-4.51$ $2.57$
$1.8$ $5.68$ $0.24$ $-4.32$ $2.61$
@@
Table 8: Survey parameters for Euclid-like luminosity Model 3, benchmark flux cut ($F_{c}=2\times 10^{-16}\mathrm{erg}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$).
[…]

### v3

@@ Appendix D MegaMapper LBG biases
$\varPhi\left(M_{\mathrm{UV}}\right)\,{\rm d}M_{\mathrm{UV}}=\left(\frac{\ln 10}{2.5}\right)\,\phi^{\star}\,10^{-0.4\,(1+\alpha)\,\left(M_{\mathrm{UV}}-M_{\mathrm{UV}}^{\star}\right)}\,\exp\left[-10^{-0.4\,\left(M_{\mathrm{UV}}-M_{\mathrm{UV}}^{\star}\right)}\right]\,{\rm d}M_{\mathrm{UV}}\;,$ (D.1)
@@
Eq. D.3 is the $K$-correction which accounts for the redshifting effect on galaxy fluxes which are measured in fixed wavelength bands. $m_{\rm UV}-m$ is negligible for the MegaMapper LBG sample, so we choose to ignore it. $D_{\rm L}(z)$ is the luminosity distance and the best-fit parameters for equation Eq. D.1 are given by Table 3 of [83][reported also here in Table 3]. This paper gives a full description of the properties of LBGs and their selection process.
Table 3: Luminosity function best-fit parameters as given by Ref. [83], which gives references for deriving the best-fit values at each redshift.
$z_{{\rm eff}}$ $M_{\rm UV}^{\star}$ $\phi^{*}/\left(10\,h^{-1}\,\mathrm{Mpc}\right)^{-3}$ $\alpha$ $m_{\rm UV}^{\star}$
$2.0$ 20.60-20.60 $9.70$ 1.60-1.60 $24.2$
$3.0$ 20.86-20.86 $5.04$ 1.78-1.78 $24.7$
$3.8$ 20.63-20.63 $9.25$ 1.57-1.57 $25.4$
$4.9$ 20.96-20.96 $3.22$ 1.60-1.60 $25.5$
$5.9$ 20.91-20.91 $1.64$ 1.87-1.87 $25.8$
@@
$n_{\mathrm{g}}\left(z,M_{\mathrm{c}}\right)=\int_{-\infty}^{M_{\mathrm{c}}(z)}\,{\rm d}M\,\varPhi(z,M),$ (D.4)
@@
$\mathcal{E}=-\frac{{\rm d}\ln n_{\mathrm{g}}}{{\rm d}\ln(1+z)}-2\,\left[1+\frac{1}{\chi\,{\cal H}}+\frac{2\,\ln 10}{5}\,\frac{{\rm d}K}{{\rm d}\ln(1+z)}\right]\,\mathcal{Q}\;.$ (D.6)
@@
$A(m)=-0.98\,(m-25)+0.11\;,\qquad B(m)=0.12\,(m-25)+0.17\;,$ (D.8)
which yields Eq. A.9 for a magnitude cut m=24.5m=24.5\,.
@@
(b) Comoving number density $n_{\rm g}$
@@
Figure 9: MegaMapper LBG luminosity function, comoving redshift distribution, and magnification and evolution biases derived from Ref. [83, 80].
@@ D.1 Luminosity derivatives
Figure 10: Comparison of the shift on the value of $f_{\mathrm{NL}}$ induced by neglecting GR corrections for MegaMapper without luminosity derivatives (solid line, as shown in Fig. 7) and with luminosity derivatives (diamond).
@@ Appendix E Survey Parameters
Table 4: Survey parameters for Roman-like luminosity Model 1 ($F_{\rm c}=1\times 10^{-16}\mathrm{erg}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$).
$z$ $V\,[\mathrm{h}^{-3}\,\mathrm{Gpc}^{3}]$ $n_{\rm g}\,[10^{-3}\,\mathrm{h}^{3}\,\mathrm{Mpc}^{-3}]$ $\mathcal{E}$ $\cal Q$}
$0.5$ $0.25$ $16.15$ 2.64-2.64 $0.81$
$0.6$ $0.32$ $13.21$ 2.82-2.82 $0.90$
$0.7$ $0.39$ $11.05$ 2.99-2.99 $0.99$
$0.8$ $0.45$ $9.39$ 3.17-3.17 $1.08$
$0.9$ $0.51$ $8.09$ 3.35-3.35 $1.17$
$1.0$ $0.56$ $7.04$ 3.52-3.52 $1.25$
$1.1$ $0.61$ $6.19$ 3.69-3.69 $1.34$
$1.2$ $0.65$ $5.48$ 3.86-3.86 $1.42$
$1.3$ $0.68$ $4.89$ 2.03-2.03 $1.50$
$1.4$ $0.71$ $4.03$ 2.19-2.19 $1.58$
$1.5$ $0.74$ $3.35$ 2.35-2.35 $1.66$
$1.6$ $0.76$ $2.81$ 2.50-2.50 $1.74$
$1.7$ $0.78$ $2.38$ 2.65-2.65 $1.81$
$1.8$ $0.79$ $2.02$ 2.80-2.80 $1.89$
$1.9$ $0.80$ $1.74$ 2.93-2.93 $1.96$
@@
Table 5: Survey parameters for Roman-like luminosity Model 3 ($F_{\rm c}=1\times 10^{-16}\mathrm{erg}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$).
$z$ $V\,[\mathrm{h}^{-3}\,\mathrm{Gpc}^{3}]$ $n_{\rm g}\,[10^{-3}\,\mathrm{h}^{3}\,\mathrm{Mpc}^{-3}]$ $\mathcal{E}$ $\mathcal{Q}$
$0.5$ $0.25$ $9.33$ 4.91-4.91 $1.07$
$0.6$ $0.32$ $7.82$ 4.71-4.71 $1.14$
$0.7$ $0.39$ $6.65$ 4.53-4.53 $1.21$
$0.8$ $0.45$ $5.70$ 4.38-4.38 $1.29$
$0.9$ $0.51$ $4.90$ 4.24-4.24 $1.36$
$1.0$ $0.56$ $4.22$ 4.12-4.12 $1.44$
$1.1$ $0.61$ $3.64$ 4.02-4.02 $1.52$
$1.2$ $0.65$ $3.13$ 3.92-3.92 $1.60$
$1.3$ $0.68$ $2.69$ 3.84-3.84 $1.69$
$1.4$ $0.71$ $2.31$ 3.76-3.76 $1.77$
$1.5$ $0.74$ $1.97$ 3.68-3.68 $1.85$
$1.6$ $0.76$ $1.69$ 3.60-3.60 $1.93$
$1.7$ $0.78$ $1.44$ 3.53-3.53 $2.01$
$1.8$ $0.79$ $1.23$ 3.45-3.45 $2.09$
$1.9$ $0.80$ $1.05$ 3.37-3.37 $2.16$
Table 6: DESI BGS survey parameters ($m_{\rm c}=20$).
$z$ $V\,[\mathrm{h}^{-3}\,\mathrm{Gpc}^{3}]$ $n_{\rm g}\,[10^{-3}\,\mathrm{h}^{3}\,\mathrm{Mpc}^{-3}]$ $\mathcal{E}$ $\cal Q$}
$0.1$ $0.11$ $38.38$ 2.66-2.66 $0.55$
$0.2$ $0.38$ $15.88$ 3.16-3.16 $0.94$
$0.3$ $0.76$ $6.42$ 3.83-3.83 $1.50$
$0.4$ $1.21$ $2.17$ 4.75-4.75 $2.31$
$0.5$ $1.69$ $0.56$ 6.03-6.03 $3.44$
$0.6$ $2.16$ $0.10$ 7.76-7.76 $4.94$
@@
Table 7: Survey parameters for Euclid-like luminosity Model 1, benchmark flux cut ($F_{\rm c}=3\times 10^{-16}\mathrm{erg}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$).
$z$ $V\,[\mathrm{h}^{-3}\,\mathrm{Gpc}^{3}]$ $n_{\rm g}\,[10^{-3}\,\mathrm{h}^{3}\,\mathrm{Mpc}^{-3}]$ $\mathcal{E}$ $\cal Q$}
$0.7$ $2.78$ $2.72$ 4.33-4.33 $1.66$
$0.8$ $3.23$ $1.98$ 4.76-4.76 $1.87$
$0.9$ $3.65$ $1.46$ 5.19-5.19 $2.09$
$1.0$ $4.03$ $1.09$ 5.62-5.62 $2.30$
$1.1$ $4.36$ $0.82$ 6.05-6.05 $2.52$
$1.2$ $4.65$ $0.62$ 6.48-6.48 $2.73$
$1.3$ $4.90$ $0.48$ 4.91-4.91 $2.94$
$1.4$ $5.12$ $0.34$ 5.33-5.33 $3.15$
$1.5$ $5.30$ $0.25$ 5.74-5.74 $3.35$
$1.6$ $5.45$ $0.18$ 6.15-6.15 $3.56$
$1.7$ $5.58$ $0.13$ 6.55-6.55 $3.76$
$1.8$ $5.68$ $0.10$ 6.94-6.94 $3.95$
$1.9$ $5.76$ $0.07$ 7.32-7.32 $4.14$
$2.0$ $5.83$ $0.06$ 7.69-7.69 $4.33$
@@
Table 8: Survey parameters for Euclid-like luminosity Model 3, benchmark flux cut ($F_{\rm c}=2\times 10^{-16}\mathrm{erg}\,\mathrm{cm}^{-2}\,\mathrm{s}^{-1}$).
$z$ $V\,[\mathrm{h}^{-3}\,\mathrm{Gpc}^{3}]$ $n_{\rm g}\,[10^{-3}\,\mathrm{h}^{3}\,\mathrm{Mpc}^{-3}]$ $\mathcal{E}$ $\cal Q$}
$0.9$ $3.65$ $1.56$ 6.13-6.13 $1.97$
$1.0$ $4.03$ $1.26$ 5.93-5.93 $2.07$
$1.1$ $4.36$ $1.02$ 5.73-5.73 $2.17$
$1.2$ $4.65$ $0.82$ 5.53-5.53 $2.26$
$1.3$ $4.90$ $0.67$ 5.32-5.32 $2.34$
$1.4$ $5.12$ $0.54$ 5.11-5.11 $2.41$
$1.5$ $5.30$ $0.44$ 4.91-4.91 $2.47$
$1.6$ $5.45$ $0.36$ 4.71-4.71 $2.52$
$1.7$ $5.58$ $0.29$ 4.51-4.51 $2.57$
$1.8$ $5.68$ $0.24$ 4.32-4.32 $2.61$
Table 9: MegaMapper LBG survey parameters ($m_{\rm c}=24.5$).
$z$ $V\,[\mathrm{h}^{-3}\,\mathrm{Gpc}^{3}]$ $n_{\rm g}\,[10^{-3}\,\mathrm{h}^{3}\,\mathrm{Mpc}^{-3}]$ $\mathcal{E}$ $\cal Q$}